An insurance representative made RM30 000 her first year and RM60 000 her seventh year. Assume that her annual salary figures form an arithmetic sequence and predict what she will make in her tenth year.


1
Expert's answer
2021-09-27T15:56:20-0400

Arithmetic progression: "a_1=30000, a_7=60000."


"a_7=a_1+d(7-1)"

"d=\dfrac{60000-30000}{6}=5000"

"a_{10}=a_1+d(10-1)"

"a_{10}=30000+5000(10-1)=75000"

She will make RM75 000 in her tenth year.